In 2017, the concentration of CO2 in the atmosphere reached 405 ppm (pCO2 = 4.05×10-4 atm). The pH of ocean water in 2017 was 8.10. If the atmospheric concentration of CO2 continues to increase at a rate of 2.3 ppm/year, what pH do you expect to observe in the oceans in the year 2037? You may assume that [HCO3-] ≈ constant because the oceans are well-buffered.
Answer #1

Concentration of CO2 in 2017 = 405ppm with a partial pressure of 4.0510-4.

It is increasing every year by the rate of 2.3 ppm/year.

So, In the year 203 , 20 years will be passed and thus the CO2 concentration will be increased by 2.320 = 46ppm.

Thus final concentration = 451 ppm. final Partial pressure = 4.0510-4. = 4.5110-4

Let the initial concentration of CO 2 in the ocean water be C1 and final be C2.

Then by Henry's law ,

C1 = KHP1 , Partial pressure of CO2 = 4.0510-4 atm.

C2 = KHP2, Partial Pressure of CO2 = 4.5110-4atm.

Thus , C2 = = 1.11 C1.

Now , pH = . Since concentration of HCO3-,is constant.

Thus , [H+]​​​​​​​finally = [H+]​​​​​​​initally

= 10-8.11.11 = 8.8110-9.

So, pH = 8.054.